OpenStudy (meilendurcer):

1 3/5 • –3/4= A. -20/24 B. -9/20 C. -1 1/5 D. - 1 9/20 the answers are negative so you know.

OpenStudy (anonymous):

\[\frac{ a }{ b } * \frac{ c }{ d } = \frac{ a*c }{ b*d }\] can you do it now ?

OpenStudy (meilendurcer):

im not sure

OpenStudy (anonymous):

\[\frac{ 3 }{5 } * \frac{ -3 }{ 4 } = \frac{ 3*-3 }{ 5*4 }\]

OpenStudy (meilendurcer):

is it B? thats what i got

OpenStudy (anonymous):

yes :)

OpenStudy (meilendurcer):

thank you